# PERC: Plan-As-Query Example Retrieval for Underrepresented Code Generation

### v1

@@ 4.1 Plan-As-Query Example Retrieval
where the in-context example for $\hat{p}_{q}$ is provided in Appendix C. As illustrated in Figure 1, Eq (4) avoids surface-level details by projecting code $c$ into plans. This is in contrast to Eq (3), which exposes the retriever to such distractions, especially when $\hat{c}$ and $c$ use different PLs.
@@ 4.2 Code Generation with Examples
When the target programming language differs from that of example code $c$, we replace $c$ with generated code $\hat{c}$ in the target language, where the LLM generates $\hat{c}$ using the in-context example shown in Appendix C:
@@ Appendix B Using C++ and Java Candidates for MultiPL-E Benchmarks
Appendix C In-Context Learning Examples
We provide concrete examples of prompts used for in-context learning in Tables 9, 10, and 11. The final prompt is created by appending the target problem to 3-shot examples in the same format as shown in the tables.
@@ Appendix C In-Context Learning Examples
Appendix D Retrieved Example and Generated Code with PERC
We provide concrete example of the few-shot examples retrieved by PERC and the resultant pseudocode and generated code on HumanEval, in which the target PL is high-resource (python), in Tables 12 and 13. We also provide examples from MultiPL-E-Lua, underrepresented target PL setting, in Tables 14 and 15.

### v2

@@ 4.1 Plan-As-Query Example Retrieval
where the in-context example for $\hat{p}_{q}$ is provided in Appendix D. As illustrated in Figure 1, Eq (4) avoids surface-level details by projecting code $c$ into plans. This is in contrast to Eq (3), which exposes the retriever to such distractions, especially when $\hat{c}$ and $c$ use different PLs.
@@ 4.2 Code Generation with Examples
When the target programming language differs from that of example code $c$, we replace $c$ with generated code $\hat{c}$ in the target language, where the LLM generates $\hat{c}$ using the in-context example shown in Appendix D:
@@ Appendix B Using C++ and Java Candidates for MultiPL-E Benchmarks
Appendix C Ablation Studies
C.1 Pseudocode as Reasoning Chain
 Benchmark MultiPL-E Lua
Reasoning Chain - Pseudocode
w/o Examples 49.94 -
Random Selection 59.38 60.68
Problem-As-Query Retrieval 60.81 60.93
CEDAR 59.69 61.61
RepoCoder 58.94 60.81
PERC 61.99 64.10
Table 9: Pass@1 scores of GPT-3.5-Turbo-16k with and without pseudocode as a reasoning chain.
C.2 Different Retrieval Methods
 Benchmark MultiPL-E Lua
Retrieval Method MPNet CodeBERT BM25
w/o Examples 49.94 49.94 49.94
Random Selection 60.68 60.68 60.68
Problem-As-Query Retrieval 60.93 58.76 60.37
CEDAR 61.61 60.99 60.25
RepoCoder 60.81 60.50 63.23
PERC 64.10 61.80 59.01
Table 10: Pass@1 scores of GPT-3.5-Turbo-16k with different retrievers.
Appendix D In-Context Learning Examples
We provide concrete examples of prompts used for in-context learning in Tables 11, 12, and 13. The final prompt is created by appending the target problem to 3-shot examples in the same format as shown in the tables.
@@ Appendix D In-Context Learning Examples
Appendix E Retrieved Example and Generated Code with PERC
We provide concrete example of the few-shot examples retrieved by PERC and the resultant pseudocode and generated code on HumanEval, in which the target PL is high-resource (python), in Tables 14 and 15. We also provide examples from MultiPL-E-Lua, underrepresented target PL setting, in Tables 16 and 17.